# Kasra Arabi, R. Teal Witter, Chinmay Hegde, Niv Cohen

### v2

@@ Abstract
Generative models have rapidly evolved to generate realistic outputs. However, their synthetic outputs increasingly challenge the clear distinction between natural and AI-generated content, necessitating robust watermarking techniques. Watermarks are typically expected to preserve the integrity of the target image, withstand removal attempts, and prevent unauthorized replication onto unrelated images. To address this need, recent methods embed persistent watermarks into images produced by diffusion models using the initial noise. Yet, to do so, they either distort the distribution of generated images or rely on searching through a long dictionary of used keys for detection.
In this paper, we propose a novel watermarking method that embeds semantic information about the generated image directly into the watermark, enabling a distortion-free watermark that can be verified without requiring a database of key patterns. Instead, the key pattern can be inferred from the semantic embedding of the image using locality-sensitive hashing. Furthermore, conditioning the watermark detection on the original image content improves robustness against forgery attacks. To demonstrate that, we consider two largely overlooked attack strategies: (i) an attacker extracting the initial noise and generating a novel image with the same pattern; (ii) an attacker inserting an unrelated (potentially harmful) object into a watermarked image, possibly while preserving the watermark. We empirically validate our method’s increased robustness to these attacks. Taken together, our results suggest that content-aware watermarks can mitigate risks arising from image-generative models. Our code is available at https://github.com/Kasraarabi/SEAL.
@@ 1 Introduction
The growing capabilities of generative models pose risks to society, including misleading public opinion, violating privacy or intellectual property, and fabricating legal evidence [14, 22, 5]. Watermarking methods aim to mitigate such risks by allowing the detection of generated contents.
Yet, many conventional watermarking techniques lack robustness against adversaries who attempt to remove them using regeneration attacks powered by recent generative models [18, 24, 9]. To address this, new watermarking techniques leveraging advances in generative models offer increased robustness against such attacks [4, 25, 27]. Namely, these methods embed a watermarking pattern in the initial noise used by a diffusion model. These patterns have been shown to be more robust against existing removal attacks.
@@
However, existing watermarks that utilize the diffusion model initial noise tend to be vulnerable to other attacks aiming to “steal” the watermark and apply it to images unrelated to the watermark owners. Some of these watermark forgery attacks can be evaded by using a distortion-free watermark - generating watermarked images from a similar distribution to the distribution of non-watermarked images; therefore exposing less information about the watermark identity. Even so, using an extensively large number of watermark identities requires maintaining a database of used noises, and might still be forgeable by other attacks.
@@
To address these challenges, we introduce SEAL -  Semantic Embedding for AI Lineage, a method that embeds watermark patterns directly tied to image semantics. Our approach enables direct watermark detection from image samples and offers the following key properties: (i) Distortion-free: As in previous works, we utilize pseudo-random hash functions to generate an initial noise that is similar to the noise used by non-watermarked models, ensuring a similar distribution of generated images. (ii) Robust to regeneration attacks: Similar to prior watermarking methods based on DDIM inversion, our approach demonstrates resilience against regeneration-based removal attempts [28]. (iii) Correlated with image semantics: The applied watermark encodes semantic information from the image. (iv) Independent of a historical database: Unlike previous methods that rely on a database of past generations, our approach embeds watermarks without requiring access to such a database, making detection possible without prior stored data.
@@
Our key insight is that we can encode semantic information about the image content in a distortion-free watermark by embedding the semantic encoding directly into the initial noise. We use projections of the user prompt embedding to seed different pseudo-random patches that compose the initial noise. We ensure the encoded embedding correlates strongly with the resulting image content, not just with the prompt, which is important since the prompt is not available during detection. At detection time, our approach identifies an image as watermarked only when the watermark pattern is both present and properly correlated with the image semantics. We describe in detail our watermarking technique in Section 3.
@@
Correlating our watermarking algorithm to the image semantics also allows us to resist forgery attacks that are challenging for many existing approaches. An attacker attempting to forge such a watermark onto unauthorized content would alter the image’s semantic embedding, breaking its correlation with the embedded pattern and rendering the watermark invalid.
@@
One mostly overlooked attack involves an adversary altering only small portions of a watermarked image while preserving the rest of its content. In such cases, the attacker can manipulate the image to be offensive, illegal, or damaging to the watermark owner’s reputation, all while the original watermark remains detectable. We term this attack the Cat Attack, as the attacker may add an object to the image (e.g., a cat) and expect the watermark to persist. We evaluate the potential damage of such attacks and demonstrate that our method uniquely provides robustness against both the Cat Attack and forgery attempts by adversaries who obtain accurate copies of our initial noise. Our experiments confirm our method’s effectiveness against these novel threats as well as previously studied attack vectors.
@@
Figure 2: Illustration of the SEAL watermarking framework for diffusion models using semantic-aware noise patterns. Watermark Generation: A textual prompt (e.g., “Beach at sunset.”) is first embedded into a semantic space. The embedding is then processed using SimHash to generate discrete keys, which are used in Encrypted Sampling to choose the initial noise $\mathbf{z}\sim\mathcal{N}(0,I)$. The watermarked noise then undergoes standard diffusion to generate the final image. Detection: The image is captioned to obtain an embedding, which is then processed by SimHash to generate a reference noise, similar to watermark generation. This noise remains correlated with the initial noise used during generation as long as the image semantics remain unchanged. The image is also processed through Inverse Diffusion to estimate the actual initial noise used during its generation. If there are insufficient matches between the reference noise and the noise obtained from inversion, the watermarking framework flags the image as non-watermarked. If a key match is found but the image is still deemed suspicious, a detailed inspection of the patches can be performed to identify local edits.
@@
  • We identify the Cat Attack, highlighting the risks of local edits to watermark owners and assessing their potential impact.
@@ 2 Related works
Recent research on image watermarking can be broadly categorized into post-processing and in-processing approaches, each offering distinct trade-offs between quality, robustness, and deployment practicality [2]. We cover here In-Processing Methods, and for Post-Processing Methods refer to Appendix 8.
In-Processing Methods.
In-processing approaches integrate watermark embedding directly within the image generation process. These methods are often used in diffusion models to achieve minimal perceptual impact. Some methods modify the generative model entirely by fine-tuning specific components, as demonstrated in Stable Signature [10]. An alternative class of techniques manipulates the initial noise of the generation process, thereby embedding the watermark without extensive model retraining. For example, Tree-Ring [25] embeds a Fourier-domain pattern into the initial noise, which can be detected through DDIM inversion [23], while RingID [7] extends this idea to support multiple keys. Other notable methods include Gaussian Shading, which produces a unique key for each watermark owner [27], PRC that leverages pseudo-random error-correcting codes for computational undetectability [12], and WIND, which employs a two-stage detection process to enables a very large number of keys [4].
@@
Recent advances in approximate nearest neighbor (ANN) search have increasingly relied on the power of Locally Sensitive Hashing (LSH) to address the challenges inherent in high-dimensional data. Originally introduced by Indyk and Motwani [13] and further refined by Gionis et al. [11], LSH employs randomized hash functions that ensure similar data points are mapped to the same bucket with high probability. Formally, for a hash family $\mathcal{H}$, the collision probability is given by
@@ 3.1 Motivation
Watermarking methods suffer from an inherent trade-off: a watermark that is harder to remove is also easier to attach to unrelated generations, compromising the reputation of the watermark owner [5]. One suggested solution to overcome this trade-off, might be maintaining a database of past generations, such that the owner could compare a seemingly watermarked image to the actual past generations. Yet, this solution is not without its problems. First, maintaining and searching a rapidly growing database, which expands with each new generation, can be challenging. Second, safeguarding the database itself may pose security risks. Finally, in various use cases, the watermark owner may not only wish to detect if an image is watermarked but also provide to a third party evidence that it is. We therefore turn to suggest a watermarking scheme that is hard to remove, hard to forge, and does not rely on maintaining a database of past generations.
Our core idea is to use a distortion-free initial noise pattern not only to indicate the origin of the image but also to encode which semantic information the image may contain. We do so in three stages (see also Figure 2): (i) Semantic Embedding – we predict a vector representing the expected semantic content in each generated image (ii) SimHash Encoding – we encode the semantic vector using a set of multi-bit hash functions (iii) Encrypted Sampling – The pseudo-random outputs of these functions are combined to produce the initial noise for the denoising process. Taken together, these steps set an initial noise that is both distortion-free with respect to standard random initialization and correlated with the semantics of the input prompt (see Section 3.3). We describe our watermarking method in detail below.
Figure 3: Effect of the Cat Attack on SEAL. (Left) A cat image is pasted onto a watermarked image at a random position and scale. (Right) Our method detects this modification by identifying elevated $\ell_{2}$ norms in affected patches.
@@ 3.2 Method
  • We do not have access to $\mathbf{v}$ at detection time; instead, we must use an approximate version $\tilde{\mathbf{v}}$ derived from the image we’re analyzing. Using $\tilde{\mathbf{v}}$, we produce an approximate version of the used initial noise $\tilde{\mathbf{z}}$.
  • Because of the randomness in the diffusion process and its inversion, we cannot reproduce exactly $\mathbf{z}$; instead, we reconstruct noise from the suspect image to obtain the inverted noise $\mathbf{z}^{\text{inv}}$.
We would ideally like for $\mathbf{z}^{\text{inv}}$ to align with $\tilde{\mathbf{z}}$ but this is not guaranteed because both differ from $\mathbf{z}$ due to the error sources mentioned above. Instead, we separate each noise vector into patches and compare them. Our method provides a high likelihood that even if some patches do not match because of the challenges discussed above, many of the patches will match as long as the suspect image is watermarked.
@@
3:  $\texttt{bits}\leftarrow\mathbf{0}$ // Initialize hash input
@@
10:  $s_{i}\leftarrow\texttt{hash}(\texttt{bits},i,\texttt{salt})$
@@
The core subroutine of our watermarking method is SimHash [6], used to generate initial noise maps correlated to a given vector (Algorithm 1). SimHash takes a vector $\mathbf{v}$ and generates an initial noise $\mathbf{z}_{i}$ for patch $i$, allowing a verifier to later determine, with some probability, whether $\mathbf{z}_{i}$ is related to $\mathbf{v}$. Namely, the semantic vector $\mathbf{v}$ is passed through a locality-sensitive hashing method that generates representations of $\mathbf{v}$ in terms of its projections in random directions.
Specifically, SimHash projects $\mathbf{v}$ onto a set of random vectors. The input to the hash function is determined by the sign of the projection, ensuring that similar vectors yield similar hash values. For $i\in\{1,\ldots,k\}$, the seed and the noise for patch $i$ are:
@@
Having repetitive patches in the initial noise may distort image generation. Therefore, we include the patch index in the hash function input to ensure that $s_{i}\neq s_{j}$ even when the input bits are identical (see Figure 8 for a visualization of what happens to the noise without the patch-dependent input). For cryptographic security, we also hash a secret salt.
Watermark Generation
Algorithm 2 Watermark Generation
1:  Input: prompt: text prompt, $n$: number of patches, $b$: number of bits, salt: secret salt
2:  Output: Watermarked image of prompt
3:  $\mathbf{z}^{\textnormal{pre}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
4:  $\mathbf{x}^{\textnormal{pre}}\leftarrow\textnormal{Diffusion}(\mathbf{z}^{\textnormal{pre}},\texttt{prompt})$
5:  $\mathbf{v}\leftarrow\textnormal{Embed}(\textnormal{Caption}(\mathbf{x}^{\textnormal{pre}}))$
6:  for $i=1,\ldots,n$ do
8:  end for
9:  return $\text{Diffusion}(\mathbf{z},\texttt{prompt})$
The first step of the generation process is to find a semantic vector describing the image that will be generated. Ideally, the semantic vector depends only on the prompt and correlates exclusively with images generated from it. Yet, in practice, predicting the image semantics based on the prompt is difficult.
Our solution begins by generating a proxy image $\mathbf{x}^{\textnormal{pre}}$. Then, we use a captioning model to achieve a text description of the generated image. The caption is embedded into a latent semantic space, resulting in a semantic vector $\mathbf{v}$, which captures the high-level semantics of the generated image by the prompt. Next, we generate the watermarked noise $\mathbf{z}$ in patches using the semantic vector $\mathbf{v}$ and SimHash. Finally, we apply diffusion to the watermarked initial noise.
Embedding Optimization. During detection, the generated image will be captioned to obtain a semantic vector $\tilde{\mathbf{v}}$. To make sure that $\mathbf{v}$ correlates to $\tilde{\mathbf{v}}$ and not to unrelated vectors, we fine-tune the embedding model to improve the similarity between the embedding of different images generated from the same prompt.
@@
1:  Input: $\tilde{\mathbf{x}}$: suspect image, $\tau:$ patch distance threshold, $n$: number of patches, $m^{\text{match}}$: match threshold, $b$: number of bits, salt: secret salt
@@
3:  $\tilde{\mathbf{v}}\leftarrow\textnormal{Embed}(\textnormal{Caption}(\tilde{\mathbf{x}}))$
4:  $\mathbf{z}^{\textnormal{inv}}\leftarrow\text{InverseDiffusion}(\tilde{\mathbf{x}})$
5:  $m\leftarrow 0$
@@
Since $\mathbf{v}$ and $\tilde{\mathbf{v}}$ may differ, $\mathbf{z}$ and $\tilde{\mathbf{z}}$ are not necessarily the same. However, by the similarity property of SimHash, $\mathbf{z}$ and $\tilde{\mathbf{z}}$ will be identical on some patches as long as $\mathbf{v}$ and $\tilde{\mathbf{v}}$ are close. On the patches $i$ where $\tilde{\mathbf{z}}_{i}=\mathbf{z}_{i}$,
@@
For such $i$, the only error stems from the diffusion and inverse diffusion processes. Empirically, we find that there is a threshold $\tau$ so that two patches have $\ell_{2}$-norm difference at most $\tau$ it is likely that they were generated from the same random seed.
Figure 4: Watermark Detection vs. Semantic Similarity. We plot the empirical probability of declaring an image as watermarked as a function of the angle between the semantic embedding used for watermark generation ($n=1024$ and $b=7$) and that of the inspected image. Our use of locally-sensitive hashing allows us to constrain the semantic embedding of the image we deem watermarked according to the initial noise used.
@@
In order to detect whether an image was initially generated with our watermark, we count the number of patches that match (i.e., their $\ell_{2}$-norm distance is at most $\tau$). If the number of matches is above a set threshold $n^{\text{match}}$ then we declare the image is watermarked. In Section 3.3, we analyze the probability of correctly identifying a watermarked image.
Tampering Detection.
In addition to the association between the watermark and the semantic embedding, edits such as object addition, removal, or modification are likely to alter the estimated initial noise in the affected image regions. This enables our watermark to provide localized information about edits that might have been made to the image. Consequently, even if the semantic embedding $\tilde{\mathbf{v}}$ aligns well with the initial embedding $\mathbf{v}$ after edits, such tampering can still be detected by identifying localized patches in the reconstructed initial noise that neither match the expected noise nor any other valid input to the hash function. Inspecting the patches one by one, the model owner may recover the $b$ input bits for each patch with an exhaustive search over the $2^{b}$ options per patch, and recover a matching initial noise.
Comparing this reconstructed noise to the inverted noise $\mathbf{z}^{\text{inv}}$ allows us to detect which patches may have been modified. The total time for this search scales as $n\cdot 2^{b}$ (which is much faster than naively searching over all $2^{(b\cdot n)}$ possible initial noise). After obtaining a per patch map (see Figure 3(b)), we may apply a spatial test as the one described in Section 10.
In any case, the local patch inspection is only required when an image is deemed watermarked by semantic similarity detection; but the watermark owner would like to have a finer understanding of the edits that might have been applied to it. This inspection is especially useful against the Cat Attack, described in Section 4.

### v3

# SEAL: Semantic Aware Image Watermarking

@@ Abstract
Generative models have rapidly evolved to generate realistic outputs. However, their synthetic outputs increasingly challenge the clear distinction between natural and AI-generated content, necessitating robust watermarking techniques to mark synthetic images. Watermarks are typically expected to preserve the integrity of the target image, withstand removal attempts, and prevent unauthorized insertion of the watermark pattern onto unrelated images. To address this need, recent methods embed persistent watermarks into images produced by diffusion models using the initial noise of the diffusion process. Yet, to do so, they either distort the distribution of generated images or require searching a large dictionary of candidate noise patterns for detection.
In this paper, we propose a novel watermarking method that embeds semantic information about the generated image into the noise pattern, enabling a distortion-free watermark that can be verified without requiring a database of key patterns. Instead, the key pattern can be inferred from the semantic embedding of the image using locality-sensitive hashing. Furthermore, conditioning the watermark detection on the original image content improves its robustness against forgery attacks. To demonstrate that, we consider two largely overlooked attack strategies: (i) an attacker extracting the initial noise and generating a novel image with the same pattern; (ii) an attacker inserting an unrelated (potentially harmful) object into a watermarked image, while preserving the watermark. We empirically validate our method’s increased robustness to these attacks. Taken together, our results suggest that content-aware watermarks can mitigate risks arising from image-generative models. Our code is available at https://github.com/Kasraarabi/SEAL.
@@ 1 Introduction
The growing capabilities of generative models pose risks to society, including misleading public opinion, violating privacy or intellectual property, and fabricating legal evidence [16, 28, 5]. Watermarking methods aim to mitigate such risks by allowing the detection of synthetically generated content.
Yet, many conventional watermarking techniques lack robustness against adversaries who attempt to remove them using regeneration attacks powered by recent generative models [12, 36, 2]. To address this, new watermarking techniques leveraging similar advances in generative models offer an increased robustness against such attacks [4, 31, 34]. Namely, these methods embed a watermarking pattern in the initial noise used by a diffusion model. These patterns have been shown to be more robust against existing removal attacks.
@@
However, existing watermarks that utilize the diffusion model initial noise tend to be vulnerable to other attacks aiming to “steal” the watermark and apply it to images unrelated to the watermark owners [32, 22, 17]. Some of these watermark forgery attacks can be evaded by using a distortion-free watermark - generating watermarked images from a similar distribution to the distribution of non-watermarked images; therefore exposing less information about the watermark identity. Even so, keeping track of a very large number of watermark identities requires maintaining a database of used noises, and might still be forgeable by other attacks [22, 17].
@@
To address these challenges, we introduce SEAL -  Semantic Embedding for AI Lineage, a method that embeds watermark patterns directly tied to image semantics. Our approach enables direct watermark detection from image samples and offers the following key properties: (i) Distortion-free: As in previous works, we utilize pseudo-random hash functions to generate an initial noise that is similar to the noise used by non-watermarked models, ensuring a similar distribution of generated images. (ii) Robust to regeneration attacks: Similar to prior watermarking methods based on DDIM inversion, our approach demonstrates resilience against regeneration-based removal attempts [36]. (iii) Correlated with image semantics: The applied watermark encodes semantic information from the image. (iv) Independent of a historical database: Our approach embeds watermarks without requiring access to a database of used noise patterns.
@@
Our key insight is that we can encode semantic information about the image content in a distortion-free watermark by embedding a semantic encoding of the generation directly into the initial noise. Namely, we may use an encoding of the requested image semantics to seed different pseudo-random patches that compose the initial noise. We ensure the encoded embedding correlates strongly with the resulting image content, and not just with the prompt, which is important since the prompt is not available during detection. At detection time, our approach identifies an image as watermarked only when the watermark pattern is both present and properly correlated with the semantics of the given image. We describe in detail our watermarking technique in Section 3.
@@
Correlating our watermarking algorithm to the image semantics also allows us to resist forgery attempts that are challenging for many existing approaches. An attacker attempting to forge our watermark onto unauthorized content would alter the image’s semantic embedding, breaking its correlation with the embedded pattern and rendering the watermark invalid.
@@
One mostly overlooked attack involves an adversary altering only small portions of a watermarked image while preserving the rest of its content. In such cases, the attacker can manipulate the image to be offensive, illegal, or damaging to the watermark owner’s reputation, all while the original watermark remains detectable. We term this attack the Cat Attack, as the attacker may add an object to the image (e.g., a cat) and expect the watermark to persist. We evaluate the potential damage of such tamperings and demonstrate that our method provides robustness against both the Cat Attack and other forgery attempts, even by adversaries who obtain accurate copies of our initial noise. Our experiments confirm our method’s effectiveness against these novel threats as well as previously studied attack vectors.
@@
Figure 2: Illustration of the SEAL watermarking framework for diffusion models using semantic-aware noise patterns. Watermark Generation: A textual prompt (e.g., “Beach at sunset.”) is first embedded into a semantic space. The embedding is then processed using SimHash to generate discrete keys, which are used in Encrypted Sampling to choose the initial noise $\mathbf{z}\sim\mathcal{N}(0,I)$. The watermarked noise then undergoes standard diffusion to generate the final image. Watermark Detection: The image is captioned to obtain an embedding, which is then processed by SimHash to generate a reference noise, similarly to watermark generation. This noise remains correlated with the initial noise used during generation as long as the image semantics remain unchanged. The initial noise is also estimated directly through Inverse Diffusion to approximate the actual initial noise used during its generation. If there are insufficient matches between the reference noise and the noise obtained from inversion, the watermarking framework flags the image as non-watermarked. If a key match is found but the image is still deemed suspicious, a detailed inspection of the patches can be performed to identify local edits.
@@
  • We investigate the Cat Attack, highlighting the risks of local edits to watermark owners and assessing their potential impact.
@@ 2 Related Works
Recent research on image watermarking can be broadly categorized into post-processing and in-processing approaches, each offering distinct trade-offs between quality, robustness, and deployment practicality [2]. We cover here In-Processing Methods, and for Post-Processing Methods refer to Section 7.
In-Processing Watermarking Methods.
In-processing approaches integrate the watermark directly within the image generation process. Some methods modify the generative model by fine-tuning specific components, as demonstrated in Stable Signature [12, 35, 26]. An alternative class of techniques manipulates the initial noise of the generation process, thereby embedding the watermark without extensive model retraining. For example, Tree-Ring [31] embeds a Fourier-domain pattern into the initial noise, which can be detected through DDIM inversion [29], while RingID [8] extends this idea to support multiple keys. Other notable methods include Gaussian Shading, which produces a unique key for each watermark owner [34], PRC that leverages pseudo-random error-correcting codes for computational undetectability [14], and WIND, which employs a two-stage detection process to enables a very large number of keys [4].
@@
Recent advances in approximate nearest neighbor (ANN) search have increasingly relied on the power of Locally Sensitive Hashing (LSH) to address the challenges of dealing with high-dimensional data. Originally introduced by Indyk and Motwani [15] and further refined by Gionis et al. [13], LSH employs randomized hash functions that ensure similar data points are mapped to the same bucket with high probability. For a hash family $\mathcal{H}$, the collision probability is given by
@@ 3.1 Motivation
Watermarking methods suffer from an inherent trade-off: a watermark that is harder to remove is also easier to attach to unrelated generations, compromising the reputation of the watermark owner [5]. One suggested solution to overcome this trade-off, might be maintaining a database of past generations, such that the owner could compare a seemingly watermarked image to the actual past generations. Yet, this solution is not without its problems. First, maintaining and searching a rapidly growing database, which expands with each new generation, can be challenging. Second, safeguarding the database itself may pose security risks. Finally, in various use cases, the watermark owner may not only wish to detect if an image is watermarked but also provide to a third party evidence that it was. We therefore turn to suggest a watermarking scheme that is hard to remove, hard to forge, and does not rely on maintaining a database of past generations.
Our core idea is to use a distortion-free initial noise pattern not only to indicate the origin of the image but also to encode which semantic information the image may contain. We do so in three stages (see also Figure 2): (i) Semantic Embedding – we obtain a vector representing the expected semantic content in each generated image (ii) SimHash Encoding – we encode the semantic vector using a set of multi-bit hash functions (iii) Encrypted Sampling – The pseudo-random outputs of the hash functions are combined to produce the initial noise for the diffusion denoising process. Taken together, these steps set an initial noise that is both distortion-free with respect to standard random initialization and correlated with the semantics of the input prompt (see Section 3.3). We describe our watermarking method in detail below.
Figure 3: Effect of the Cat Attack on SEAL. (Left) A cat image is pasted onto a watermarked image at a random position and scale. (Right) Our method detects this modification by identifying elevated $\ell_{2}$ norms in affected patches. Note that the displayed norms are rounded to the nearest integer.
@@ 3.2 Method
  • We do not have access to $\mathbf{v}$ at detection time; instead, we must use an approximate version $\tilde{\mathbf{v}}$ derived from the image we are analyzing. Using $\tilde{\mathbf{v}}$, we produce an approximate version of the used initial noise $\tilde{\mathbf{z}}$.
  • Because of the randomness in the diffusion process and its inversion, we cannot estimate $\mathbf{z}$ accurately; instead, we get an approximation of the inverted noise $\mathbf{z}^{\text{inv}}$.
Ideally, a watemarked image would yield a perfect match between the noisese derived from the image semantics $\mathbf{z}^{\text{inv}}$ and the inverted $\tilde{\mathbf{z}}$ but this is not guaranteed because both differ from $\mathbf{z}$ due to the error sources mentioned above. Yet, we can mitigate this by independently embedding the image semantics across multiple patches. Therefore, our method provides a high likelihood that even if some patches do not match because of the challenges discussed above, many of the patches will match as long as the suspect image is watermarked.
Algorithm 1 Watermark Generation
1:Input: prompt: text prompt, $n$: number of patches, $b$: number of bits per patch, salt: secret salt
2:Output: Watermarked image of prompt
[…]
6:for $i=1,\ldots,n$ do
8:end for
9:return $\text{Diffusion}(\mathbf{z},\texttt{prompt})$
@@
[…]
@@
[…]
@@
Watermark Generation
The first step of the generation process is to find a semantic vector $\mathbf{v}$ describing the image that will be generated. Ideally, the semantic vector depends only on the prompt and correlates exclusively with images generated from it. Yet, in practice, predicting the final image semantics based on the user prompt is difficult.
To approximate the generated image semantics, our solution begins by generating a proxy image $\mathbf{x}^{\textnormal{pre}}$. We first caption the image using BLIP-2 model [19]. Then, the caption is embedded into a latent semantic space using the Paraphrase Mpnet Base V2 model [24], resulting in a semantic vector $\mathbf{v}$ which captures the high-level semantics of the generated image by the prompt (a similar concept also explored in the concurrent work of SWIFT [10]).
Semantic Embedding Optimization. During detection, the generated image will be captioned to obtain a semantic vector $\tilde{\mathbf{v}}$, approximating the semantic vector $\mathbf{v}$ that was used to seed the random noise. To ensure a similarity between $\mathbf{v}$ and $\tilde{\mathbf{v}}$, it is not enough only to generate the proxy image $\mathbf{x}^{\textnormal{pre}}$ with the same prompt (a qualitative comparison of the captioned proxy image and the final generated image can be found at Figure 9). Therefore, to encourage the embedding $\mathbf{v}$ to correlate to $\tilde{\mathbf{v}}$ and not to unrelated vectors, we fine-tuned the embedding model using 10k pairs of related captions, leading to additional improvements (Figure 5). The full implementation details of the fine-tuning process can be found in Section 9.4.
After obtaining the desired semantic embedding, we generate the watermarked noise $\mathbf{z}$ using the semantic vector ${\mathbf{v}}$ and the SimHash algorithm described below. Finally, we will use the diffusion mode to generate the image with the watermarked initial noise. The generation algorithm is summarized in Algorithm 1.
@@
The core subroutine of our watermarking method is SimHash [7], used to generate initial noise patches correlated to a given vector (Algorithm 2). SimHash takes a vector $\mathbf{v}$ and generates an initial noise $\mathbf{z}_{i}$ for patch $i$, allowing a verifier to later determine whether $\mathbf{z}_{i}$ is related to $\mathbf{v}$. Namely, the semantic vector $\mathbf{v}$ is passed through a locality-sensitive hashing method that generates representations of $\mathbf{v}$ in terms of its projections onto random directions.
Specifically, SimHash projects $\mathbf{v}$ onto a set of random vectors for each patch of the initial noise map. It uses $b$ projection vectors for each of the $k$ noise patches. Each noise patch is generated using a seed determined by the sign of the projection of the semantic vector onto each of the $b$ projection directions. For $i\in\{1,\ldots,k\}$, the seed and the noise for patch $i$ are:
@@
This ensures that similar semantic vectors would yield similar hash values.
Yet, having repetitive bit inputs ($s_{i}$) may result in repetitive patches in the initial noise, and therefre may distort image generation. Therefore, we include the patch index in the hash function input to ensure that $s_{i}\neq s_{j}$ even when the input bits are identical (see Figure 8 for generation samples in the case of repetitive noise patches). For cryptographic security, we also use a user-specific secret salt.
@@
1:Input: $\tilde{\mathbf{x}}$: suspect image, $\tau:$ patch distance threshold, $n$: number of patches, $m^{\text{match}}$: match threshold, $b$: number of bits per patch, salt: secret salt
@@
[…]
@@
Since $\mathbf{v}$ and $\tilde{\mathbf{v}}$ may differ, the originally used noise $\mathbf{z}$ and $\tilde{\mathbf{z}}$ are not necessarily the same. However, by the similarity property of SimHash, $\mathbf{z}$ and $\tilde{\mathbf{z}}$ will be identical on some patches with very high probability as long as $\mathbf{v}$ and $\tilde{\mathbf{v}}$ are close. On any patch number $i$ where the SimHash patch match ($\tilde{\mathbf{z}}_{i}=\mathbf{z}_{i}$), we get:
@@
[…]
Angle $\theta(\mathbf{v},\tilde{\mathbf{v}})$ Detection Probability
$65^{\circ}$ $8.55\times 10^{-4}$
$60^{\circ}$ 0.0530.0530.053
$55^{\circ}$ 0.5510.5510.551
$50^{\circ}$ 0.9980.9980.998
$45^{\circ}$ 1.0001.0001.000
Figure 4: Watermark Detection vs. Semantic Similarity. We plot the empirical probability of detecting a watermark as a function of the angle between the semantic embedding used for watermark generation and that of the inspected image ($n=1024$, $b=7$). The table shows the analytical detection probabilities at key angles calculated by Lemma 3.2, illustrating how sharply SimHash distinguishes semantically related images from unrelated ones.
@@
Finally, in order to detect whether an image was initially generated with our watermark, we count the number of patches that match (i.e., their $\ell_{2}$-norm distance is below a threshold $\tau$). If the number of matches is above a set threshold $n^{\text{match}}$ then we declare the image is watermarked. In Section 3.3, we analyze the probability of correctly identifying a watermarked image.
Tampering Detection with a Spatial Test.
In addition to the association between the watermark and the semantic embedding, edits such as object addition, removal, or modification are likely to alter the estimated initial noise in the affected image regions. This enables our watermark to provide localized information about edits that might have been made to the image. Consequently, even when the semantic embedding of the image $\tilde{\mathbf{v}}$ aligns well with the initial embedding used to seed the noise $\mathbf{v}$, such tampering edits can still be detected by identifying localized patches in the reconstructed initial noise that neither match the expected noise nor any other valid input to the hash function.
To detect such cases, we may inspect the noise patches one by one. Given the model owner’s private information, we may recover the $b$ input bits used to seed each patch with an exhaustive search over the $2^{b}$ options per patch, and recover a matching initial noise. Comparing this reconstructed noise to the inverted noise $\mathbf{z}^{\text{inv}}$ allows us to detect which patches may have been modified. The total time for this search scales as $n\cdot 2^{b}$ (which is much faster than naively searching over all $2^{(b\cdot n)}$ possible initial noise). After obtaining a per patch noise-matching map (see Figure 3(b)), we may apply a spatial test as the one described in Section 9.2 to detect tampering attempts. In any case, the local patch inspection is only required when an image is deemed watermarked by semantic similarity detection; but the watermark owner would like to have a finer understanding of the edits that might have been applied to it. This inspection is especially useful against the Cat Attack, described in Section 4.